OpenStudy (anonymous):

Help, number bases,321x=57 base 10

OpenStudy (anonymous):

convert 321 from base x to 10 i.e. 3x^2+2x+1=57

OpenStudy (anonymous):

how then do you work out the quadratic equation?

OpenStudy (anonymous):

use the formula

OpenStudy (anonymous):

 this root is equal to what. calc. pls.

OpenStudy (apoorvk):

you mean \(\sqrt{-671}\) ??

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

Okay, \(\sqrt{-671} = \sqrt{671} \times \sqrt{-1}\), and since \(\sqrt{-1}\) is known as \(\large i\), or 'iota' (imaginary number) so, this whole thing equals \(\sqrt{671}i\).

OpenStudy (anonymous):

the roots are 4 and -4.67. since the base cannot be a negative number, x=4